OpenStudy (anonymous):

x^3+512=0 please i need to see steps

OpenStudy (mathstudent55):

Is 512 a perfect cube?

OpenStudy (mathstudent55):

\(a^3 - b^3 = (a - b)(b^2 + ab + b^2) \)

OpenStudy (anonymous):

okay

OpenStudy (mathstudent55):

512 = 8^3

OpenStudy (mathstudent55):

Sorry, you need the sum of two cubes, not the difference. \(\large a^3 + b^3 = (a + b)(a^2 - ab + b^2)\)

OpenStudy (mathstudent55):

You have \(\large x^3 + 8^3 = 0\) Now just factor following the sum of two cubes pattern just above.

OpenStudy (mathstudent55):

\(\large (x + 8)(x^2 - 8x + 64) = 0 \) \(\large x + 8 = 0\) or \(\large x^2 - 8x + 64 = 0 \)

OpenStudy (mathstudent55):

The left equation is simple. Use the quadratic formula on the right equation.

OpenStudy (anonymous):

looking for how to breakdown -192

OpenStudy (anonymous):

on the root

OpenStudy (mathstudent55):

\(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \) \(x = \dfrac{-(-8) \pm \sqrt{(-8)^2 - 4(1)(64)}}{2(1)} \)

OpenStudy (mathstudent55):

\(x = \dfrac{8 \pm \sqrt{64 - 256}}{2} \) \(x = \dfrac{8 \pm \sqrt{-192}}{2} \)

OpenStudy (mathstudent55):